# Space Surveillance with High-Frequency Radar

### v1

@@ II System Description
The HFLOS transmit system consists of multiple elements, with the total transmit power arising from freespace combination from a transmit array. This allows for a modular design and deploymen2 t of the waveform generators, power amplifiers, and transmit antenna for each transmit channel. HF wavelengths ensure a large beam for wide-area illumination. This coverage can be broadened with physical pointing or phase adjusting (beam-spreading) the transmit array to form an instantaneous wide field of regard.
@@ IV Correction for Ionospheric Effects
Two different methods are used to calculate the correction to the range rate (Doppler). The first, referred to as the Phase Path method, requires the range rate to be calculated using the phase path. The Doppler shift, $\delta f$, is [69]
$\delta f=-\frac{f}{c}\frac{dP}{dt},$ (14)
where $f$ is the radio wave frequency and $c$ the speed of light in vacuo. As range rate is related to Doppler shift by $-\frac{c\delta f}{2f}$, we have the following equation for the corrected range rate:
$\frac{dR_{c}}{dt}=\frac{1}{2}\frac{dP}{dt}.$ (15)
However, random fluctuations in the radar range measurements are replicated in the phase path. Calculating the numerical derivative can therefore lead to uncertainties becoming unsuitably large. A second method, referred to as the Range Rate method was developed to mitigate this. The Range Rate method requires the applicability of the quasi-longitudinal approximation, requires the ionosphere to be free of spatial or temporal gradients over the RSO observation, and assumes that the contribution to the range correction by ray-bending is small compared to that of group retardation. A discussion on the derivation of this method is provided in [46].
The corrected range rate is calculated in terms of the measured range rate, $\frac{dR_{m}}{dt}$, and the partial time derivative of the range correction, $\delta R$, which is in turn the difference of the corrected and measured range ($\delta R=R_{c}-R_{m}$):
$\frac{dR_{c}}{dt}=\frac{dR_{m}}{dt}+2\frac{\partial(\delta R)}{\partial t}.$ (16)

### v2

@@ II System Description
The HFLOS transmit system consists of multiple elements, with the total transmit power arising from freespace combination from a transmit array. This allows for a modular design and deployment of the waveform generators, power amplifiers, and transmit antenna for each transmit channel. HF wavelengths ensure a large beam for wide-area illumination. This coverage can be broadened with physical pointing or phase adjusting (beam-spreading) the transmit array to form an instantaneous wide field of regard.
@@ IV Correction for Ionospheric Effects
Two different methods are used to calculate the correction to the range rate (Doppler). The first, referred to as the Raytracing method, calculates the corrected range rate by taking the numerical derivative of the corrected range, calculated above:
$RR_{c}=\frac{dR_{c}}{dt},$ (14)
where $RR_{c}$ is the corrected range rate and $R_{c}$ is the raytraced corrected range.
However, random fluctuations in the radar range measurements are replicated in the corrected range. Calculating the numerical derivative can therefore lead to uncertainties becoming unsuitably large. A second method, referred to as the Phase-Path Correction method was developed to mitigate this. Consider the measured Doppler shift, $\delta f$ [69]:
$\delta f=-2\frac{f}{c}\frac{dP}{dt},$ (15)
where $f$ is the radio wave frequency, $c$ the speed of light in vacuo, $P$ is the phase path, and the factor of 2 accounts for propagation both from the radar to the RSO and back. As range rate is related to Doppler shift by $-\frac{c\delta f}{2f}$, we have the following equation for the measured range rate:
$RR_{m}=\frac{dP}{dt}.$ (16)
Since the phase path is also calculated by PHaRLAP during the ray tracing process, the phase path correction, $\delta P=R_{c}-P$ can be calculated independently of the range rate measurement. In this expression, taking the difference between $R_{c}$ and $P$ effectively subtracts any errors which were replicated from the measured range. We can now take the derivative:
$\frac{d(\delta P)}{dt}=\frac{dR_{c}}{dt}-\frac{dP}{dt}.$ (17)
Substituting in the expressions in Equations 14 and 16, we have a second, independent method of calculating the corrected range rate:
$RR_{c}=RR_{m}+\frac{d(\delta P)}{dt}.$ (18)